by what number should (5/6)^-3 be divided so the quiotent be 6^3*5^-2

answer with steps asap
1
Expert's answer
2020-12-16T19:21:47-0500

"\frac{5}{6}^{-3} =" "\frac{216}{125}"

"6^3 \times 5^{-2} =" "= \frac{216}{25}"

"\frac{216}{125} \times x = \frac{216}{25}"

"x = \frac{216}{25} \times \frac{125}{216}"

"x = 5"

Hence the number should be divided by "\frac{1}{5}"